OpenStudy (anonymous):

I need help with this equation:

OpenStudy (anonymous):

\[\left(\begin{matrix}a ^{2}+1 \\ a ^{2}-1\end{matrix}\right)-\left(\begin{matrix}2 \\ a+1\end{matrix}\right)=\left(\begin{matrix}a \\ a+1\end{matrix}\right)\]

OpenStudy (anonymous):

sweet jesus where would this pop up naturally lol

myininaya (myininaya):

are those suppose to fractions?

OpenStudy (anonymous):

yes their fractions

OpenStudy (anonymous):

i think its nCr \[\frac{n!}{r!(n-r)!}\]

myininaya (myininaya):

\[\frac{a^2+1}{a^2-1}-\frac{2}{a+1}=\frac{a}{a+1}\]

myininaya (myininaya):

lol joe wouldnt that be a cool problem

OpenStudy (anonymous):

i was already on it lolol

myininaya (myininaya):

first of all a^2-1 is factorable right? what are its factors?

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

the answer is 1.

myininaya (myininaya):

you have to know how to factor before you can do a problem like this? a^2-1 is the difference of two squares you should have a formula that you use to factor it a^2-b^2=(a-b)(a+b) so what are the factors of a^2-1?